# Adaptive Manipulation using Behavior Trees

### v2

@@ VII Experiments
To demonstrate the modularity offered by the adaptive BT, we also consider a walk-and-manipulate task where the robot must navigate to, tighten, and navigate away from the needle valve. The adaptive BT is inserted as a subtree within a larger BT that also commands navigation to and from the device (using the standard ANYmal C navigation and locomotion software). Fig. 5a-5c illustrates this experiment.
Figure 4: An annotated timeline of tightening the needle valve using the adaptive BT with both available strategies, with plots of the effort in the wrist joint during each manipulation attempt A, B, C and D. Note the effort spikes during attempts C and D; the first indicates that the low-torque strategy is no longer feasible, while the second indicates that the device is sufficiently tightened. For clarity, only the lowest active effort limit is shown in each plot.
@@
Figure 5: Snapshots of experiments performed. (5a) Navigating to the needle valve. (5b) Tightening the needle valve. (5c) Navigating away from the needle valve. (5d) Pushing the e-stop button until fully pressed. (5e) Twisting the e-stop button until released. (5f) Tightening the globe valve.
E-Stop Button. The robot manipulates an IDEC HW1X-BV411-R emergency stop[46]; it pushes the button until fully pressed and twists it until released. One strategy was provided for each task; Fig. 5d and 5e depict these strategies. A couple thin foam strips were placed on the Unitree Z1’s gripper to allow it to tightly grasp the button for twisting.
Globe Valve. The robot twists a Dikkan DN40 PN16 globe valve[47] from unknown starting tightness until tight. One strategy was provided for this task; Fig. 5f depicts this strategy. The valve’s diaphragm and O-ring were removed to make it loose enough to be twisted by the Unitree Z1, and some thin foam strips were placed on the handle to allow the Z1’s gripper to tightly grasp the handle.
@@ VIII Results and Discussion
An accompanying video for the experiments performed is provided with this work [48].
Each experiment was repeated for 10 trials with the device handle pose and recorded effort data reset for each trial. Strategy effort limits were set to the Unitree Z1’s maximum safe operating efforts scaled down by a small safety factor, except for the low-torque strategy which used a larger safety factor to avoid the base becoming unstable from the arm’s wrench during heavy twisting. Task success was verified by checking the device after the robot had exited the task. Table I records, for each experiment, the cumulative success rate after each attempt, ignoring re-attempts triggered to change strategy and re-grasping when device handle pose limits were exceeded, as well as times taken for task completion. A maximum of 5 attempts was permitted for each trial. Fig. 4 provides an annotated timeline of tightening the needle valve when having access to both associated strategies.
TABLE I: Results for All Manipulation Tasks Using the Adaptive BT
Device Experiment Cumulative Success Rate (%) After Attempt Number Time Taken (s)
@@
Needle Valve Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 135 (Fail) 136 (Fail)
@@
Both Strategies (Adaptive) 80 100 Done Done Done 175 182
@@
Globe Valve Tighten 70 100 Done Done Done 95 107
@@
Table I shows that the robot achieved a 100% success rate within the permitted number of attempts for all but one experiment when using the adaptive BT, despite having no knowledge of the valves’ initial tightness, or how much they needed to be twisted to be tightened. The robot never suffered catastrophic failure since the condition monitor preempted failure before it occurred. When only having access to the low-torque strategy, the robot aborted tightening the needle valve after the measured joint effort exceeded the strategy effort limit for the arm’s wrist joint. This demonstrates the safety enabled by the adaptive BT, since the robot was able to react and safely exit the task before overexerting itself, and would not try again because the manipulation data suggested that no feasible strategies were available. However, the utility of having access to the low-torque strategy is made clear when comparing the task completion times for using the high-torque strategy only and adaptively using both strategies. Since the low-torque strategy is approximately 5 times faster while twisting, it is optimal for the robot to use it for most of the manipulation sequence, but when the robot experiences a sudden effort spike as the valve begins to tighten, it quickly adapts and uses the high-torque strategy to fully tighten the valve; this resulted in an overall task speedup of 46% on average. We note that this is similar to behaviour displayed by humans, twisting quickly when the valve is loose but changing grip near the end to be able to apply more wrench without hurting oneself. The robot was also able to safely twist the valves for an arbitrary number of rotations without the danger of overextending itself. This was done by periodically re-grasping the valves using the condition of exceeding safe device pose limits; once triggered, the robot reattempted the task and re-grasped the device in a more favorable joint configuration for further twisting.
Attempting these tasks with the same robot, motion planner, controllers and strategy subtree but without the condition monitor (instead executing a plan made prior to manipulation without adapting to effort observations during execution) resulted in a 0% success rate for all tasks, and effort spikes frequently resulted in catastrophic failure with the Z1 shutting down from overheating or blowing a fuse. This highlights the necessity of adaptive behavior for these tasks.
@@ IX Conclusion and Future Work
In this work we proposed the adaptive BT, a flexible and highly modular BT design that enables a robot to adapt to an uncertain environment by pairing reactive condition monitoring with adaptive strategy selection. Our test robot used the adaptive BT to safely and reliably complete complex effort-intensive manipulation tasks commonly found in industrial settings, despite using out-the-box controllers. The adaptive BT can be readily added to existing BT implementations without major changes, enhancing robustness while remaining interpretable and without significantly changing the system architecture. Future work will investigate algorithmic and learning-based synthesis of new strategies and conditions for the adaptive BT to enable a robot to generalize to new tasks, including synthesis from more semantic, conceptual descriptions of each strategy, or autonomous discovery of new strategies. We also plan to benchmark against other TAMP methods as well as run experiments beyond the lab.
Acknowledgments
The authors thank Emir Kocak for the implementation and deployment of the motion planner and control interface.
[Actuated Devices in Simulation]
Preliminary testing was done in simulation using Gazebo [49], using simulated devices that are highly accurate to their real-world equivalents, allowing for testing to closely reflect testing on hardware. They can be accessed at [50].

### v3

@@ VII Case Studies
Needle Valve. The robot twists a Tameson NLS-012 needle valve from unknown starting tightness until tight, and loosens it to a user-specified angle, with access to the two strategies depicted by Fig. 1. To demonstrate how the adaptive BT can be readily placed within larger BT structures, we also consider a walk-and-manipulate task where the robot navigates to, tightens, and navigates away from the needle valve (with access to both manipulation strategies, but the required applied torque for tightness reduced so as to be achievable by the ‘low-torque’ strategy). The adaptive BT is inserted as a subtree within a larger BT that also commands navigation to and from the device (using the standard ANYmal navigation and locomotion software).
Globe Valve. The robot twists a Dikkan DN40 PN16 globe valve from unknown starting tightness until tight, with access to a single strategy depicted by Fig. 5b.
E-Stop Button. The robot manipulates an IDEC HW1X-BV411-R emergency stop; it pushes the button until fully pressed, with access to a single strategy depicted by Fig. 5c, and twists it until released, with access to a single strategy very similar to that depicted by Fig. 1a.
Groceries Packing. The robot packs a plastic bottle, (full) juice box, (real) orange and (plastic) banana into a basket, in a user-specified order, as depicted by Fig. 5d. The grasping stage uses a single strategy, with subtree structures depicted by Fig. 4b, determining firm grasping by checking gripper joint effort. The packing stage has access to a continuous set of strategies; all strategies have the same subtree structures depicted by Fig. 4c, but the speed of the motion trajectories is parameterized by a speed factor (i.e. a strategy) $a\in\mathcal{A}_{\mathrm{safe}}$, where $\mathcal{A}_{\mathrm{safe}}$ is determined by the weight of the object (lower maximum safe speeds for heavier objects, which may slip out of grasp when the motions are too fast). The adaptive BTs for the two stages are placed in sequence such that the robot returns to the first stage if the second stage is preempted.
@@
Figure 5: Snapshots from experiments. (5a) Walking to the needle valve (as part of the walk-and-manipulate task). (5b) Tightening the globe valve. (5c) Pushing the e-stop button until fully pressed. (5d) Groceries packing. Accompanying video with experiments: https://youtu.be/im7OCkEiQsU.
Baseline. For all tasks, we use backward chaining with stacked Fallbacks as the baseline BT structure [15]. This is the standard BT design used when the robot must choose between different actions with robustness to failure. For fair comparison to the adaptive BT, we manually arranged the (nominal) ‘faster’ strategies before ‘slower’ ones in the baseline’s Fallbacks such that the robot tries them first, with a reminder that the adaptive BT automates (and optimizes) the strategy selection process. Since the baseline cannot select from a continuous set of strategies, it can only (safely) use the lowest speed factor when grocery packing.
@@ VIII Results and Discussion
To evaluate the adaptive BT, we performed a series of ablations to assess how each component impacts performance, with comparison to the baseline where relevant. For our discussion, we refer to the results in Table I, which records, for 10 trials of each experiment, the cumulative success rate after each attempt (ignoring re-attempts triggered to change strategy and re-grasping when device handle pose limits were exceeded) and time taken for task completion. A maximum of 5 attempts was permitted for each trial, task success was verified by checking the scene after the robot had exited the task, and the initial device states were reset to be the same for each trial. Strategy effort limits were set to the Z1’s maximum safe operating efforts scaled down by a safety factor, except for the needle valve low-torque strategy, which used a larger safety factor to avoid the quadruped becoming unstable from induced wrench during heavy twisting.
TABLE I: Experiment Results for All Tasks
Task Experiment Cumulative Success Rate (%) After Attempt Number Time Taken (s)
@@
Ablation A: RFP Enabled (SAT + LPE Disabled, i.e. Equivalent to Baseline)
Needle Valve (Tighten) Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 135 (Fail) 136 (Fail)
@@
Needle Valve (Loosen) Low-Torque Strategy Only 0 Aborted Aborted Aborted Aborted 58 (Fail) 58 (Fail)
High-Torque Strategy Only 70 90 100 Done Done 228 247
Globe Valve Tighten 70 100 Done Done Done 95 107
@@
Groceries Packing Lowest Speed Only 60 70 90 100 Done 261 269
Ablation B: RFP + SAT Enabled (LPE Disabled)
Needle Valve (Tighten) Both Strategies (Baseline) 70 100 Done Done Done 175 186
Both Strategies (Adaptive BT) 80 100 Done Done Done 175 182
Needle Valve (Loosen) Both Strategies (Baseline) 60 90 100 Done Done 259 283
Both Strategies (Adaptive BT) 80 100 Done Done Done 204 212
Groceries Packing Lowest Speed Only (Baseline) 60 70 90 100 Done 261 269
Weight-Based Speed (Adaptive BT) 50 70 80 100 Done 196 206
Ablation C: RFP + SAT + LPE Enabled (Complete Adaptive BT)
Needle Valve (Tighten) Both Strategies (Baseline) 80 100 Done Done Done 94 126
Both Strategies (Adaptive BT) 90 90 100 Done Done 58 118
Needle Valve (Loosen) Both Strategies (Baseline) 60 90 100 Done Done 259 283
Both Strategies (Adaptive BT) 80 100 Done Done Done 172 182
@@
Reactive Failure Preemption (RFP). For Ablation A, the robot only has access to one strategy (preventing strategy adaptation) and manipulation data is reset between trials (preventing learning from past experience). We see a 100% success rate within the permitted number of attempts for all but one experiment when using the adaptive BT, despite having no knowledge of the valves’ initial tightness, or how much they needed to be twisted to be tightened. The robot never suffered catastrophic failure since the condition monitor preempted failure before it occurred. When using the low-torque strategy, the robot aborted the needle valve tasks after exceeding the safe effort limit for the arm’s wrist joint. This demonstrates the safety enabled by the adaptive BT, since the robot was able to react and safely exit the task before overexerting itself. The robot was also able to safely twist the valves for an arbitrary number of rotations without the danger of overextending itself, by periodically re-grasping the valves using the condition of exceeding safe device pose limits; once triggered, the robot re-attempted the task and re-grasped the device in a more favorable joint configuration for further twisting. The robot was also able to firmly grasp the different-shaped fragile objects without damaging them by reacting to spikes in gripper effort.
Under this ablation, the adaptive BT produces equivalent behaviors to the baseline. However, the adaptive BT enables easier scalability of task composition over the baseline, since the adaptive BT provides an abstracted, modular interface for appending and removing subtrees for strategies and conditions to the strategy database (see Fig. 3), while editing the baseline backward chained BT structure can become unwieldy as the number of strategies and conditions grows.
Strategy Adaptation (SAT). For Ablation B, the robot has access to all strategies but manipulation data is still reset between trials (preventing learning from past experience). We see substantial average task speedups for needle valve tightening (46%) and loosening (14%) as well as groceries packing (23%) when using the adaptive BT compared to Ablation A; the cause for this speedup is that the adaptive BT selects faster strategies when feasible and switches to slower, ‘safer’ strategies when necessary.
For needle valve tightening, since the low-torque strategy is approximately 5 times faster while twisting, it is optimal for the robot to use it for most of the manipulation sequence, but when the robot experiences a sudden effort spike as the valve begins to tighten, it quickly adapts and uses the high-torque strategy to fully tighten the valve; Fig. 6 provides an annotated timeline. We note that this is similar to behavior displayed by humans, twisting quickly when the valve is loose but changing grip near the end to be able to apply more wrench without hurting oneself. For this particular task the baseline produces the same behavior, leading to roughly the same average time taken; this is because the optimal behavior for this task (start low-torque, switch to high-torque at the end) can also be achieved using a simple Fallback. However in general this is not the case, as demonstrated by the needle valve loosening task, wherein the adaptive BT achieves a 25% average speedup over the baseline. For this task, both the adaptive BT and baseline first (unsuccessfully) try the low-torque strategy (which cannot safely apply enough torque to loosen the valve) and then switch to the high-torque strategy for the first successful twist, after which the valve is loose enough to be loosened further using the low-torque strategy. While the logic of the Fallback means that the baseline will keep using the high-torque strategy for the remainder of the task, the adaptive BT optimizes over feasible strategies at each re-attempt, and thus chooses the (faster) low-torque strategy for the remainder of the task.
Figure 6: An annotated timeline of tightening the needle valve using the adaptive BT with both available strategies, with plots of the effort in the wrist joint during each manipulation attempt A, B, C (which all use the ‘low-torque’ strategy) and D (which uses the ‘high-torque’ strategy). The adaptive BT initially selects the low-torque strategy to twist quickly while the valve is loose but switches to the high-torque strategy near the end to apply more wrench and finish tightening. Note the effort spikes during attempts C and D; the first indicates that the low-torque strategy is no longer feasible, while the second indicates that the device is sufficiently tightened. For clarity, only the lowest active effort limit is shown in each plot.
Learning from Past Experience (LPE). For Ablation C, the robot has access to all strategies as well as the manipulation data recorded from Ablations A and B; note that the baseline has no mechanism for learning from past manipulation data. For needle valve loosening, the adaptive BT achieves an even greater average speedup (36%) over the baseline; this is because past manipulation data suggests that the valve will initially be too tight to be loosened with the low-torque strategy, so the robot has learned to start with the high-torque strategy (while in Ablation B, due to lack of knowledge of the valve’s initial tightness, it would initially try low-torque then immediately switch to high-torque).
For needle valve tightening, we see no change from Ablation B if the robot has no initial estimate of the required rotation until tight, i.e. task state $s$. Note that as we vary initial $s$, in some cases it is faster to start with the high-torque strategy, since it avoids the overhead of switching from low-torque to high-torque to finish tightening. If we now provide the adaptive BT with an accurate initial estimate for $s$, it always starts with the strategy that minimizes time taken to tighten the valve according to past experience (unlike Ablation B, which always starts with low-torque). This is illustrated by Fig. 7, which plots minimum time taken to tighten the needle valve for 5 different known required rotations, with the adaptive BT achieving up to 34% speedup over the baseline (at 0.5 radians). For Ablation C’s results for this task in Table I, the 10 trials are instead 2 trials each over these 5 different known required rotations; here the adaptive BT achieves 6.3% average speedup over the baseline.
In practice, initial $s$ could be a random variable with a known distribution; for this case, the mean time taken over this distribution would be the appropriate $t_{\mathrm{est}}$ to use for strategy selection. If the distribution is unknown, it could be estimated empirically over many manipulations.
Figure 7: Minimum time to tighten the needle valve for different (known) required rotations until tight. The adaptive BT starts with the strategy that minimizes time taken to tighten the valve according to past experience.
@@ IX Conclusion and Future Work
In this work we proposed the adaptive BT, a flexible and highly modular BT design that enables a robot to adapt to environment uncertainties by pairing reactive condition monitoring with adaptive strategy selection, and learning from past experience to predict the optimal strategy given the task state. Our test robot used the adaptive BT to safely and robustly complete a diverse range of complex effort-intensive manipulation tasks commonly found in industrial and domestic settings, despite using out-the-box controllers, and with reduced time taken over the baseline BT design for all experiments where multiple strategies were available. The adaptive BT can be readily added to existing BT implementations without major changes, enhancing performance while remaining interpretable and without significantly changing the system architecture. Future work will investigate algorithmic and learning-based synthesis of new strategies and conditions for the adaptive BT to enable a robot to generalize to new tasks, including synthesis from more semantic, conceptual descriptions of each strategy, or autonomous discovery of new strategies.
Appendix A Computing The Feasible Set of Strategies
When manipulating articulated devices, all feasible strategies for task state $s$ are such that the robot arm joint efforts are not exceeded when using those strategies in the window $\{s^{\prime}:s_{L}\leq s^{\prime}-s\leq s_{U}\}$ around $s$. Thus, $\mathcal{A}_{\mathrm{safe}}$ is given by
[…]
$\forall i\in[1..N_{\mathrm{DOF}}],s_{L}\leq s^{\prime}-s\leq s_{U}\},$
where $T_{\mathrm{est},i}(a,s)$ are the estimated efforts for each arm joint $i$ at task state $s$ using strategy $a$, recorded from previous manipulation attempts, and $T_{\mathrm{lim},i}(a)$ are the maximum safe efforts for each joint under strategy $a$. $T_{\mathrm{est},i}$ are recorded from the joint efforts during each manipulation attempt. For valve manipulation, the task state $s$ is the total rotation angle $\theta$ away from being tight, with initial estimate 0 radians for loosening tasks. For tightening tasks, initial $\theta$ is estimated to be a large default value if unknown, but once the valve is tight, all $\theta$ values in the new batch of manipulation data are shifted such that the final entry is 0 radians. For button manipulation, the task has a (known) binary on/off state.
When grocery packing, the speed factor is computed from the additional effort that the weight of the object induces in the arm joints. Accordingly, $\mathcal{A}_{\mathrm{safe}}$ is given by
$\mathcal{A}_{\mathrm{safe}}(s)=\{a\in\mathbb{R}^{+}:a_{L}\leq a\leq a_{H}(s)\}$
where $\text{where }a_{H}(s)=\min\bigl{\{}a_{L},1-\sum_{i}\alpha_{i}|T_{\mathrm{obj},i}(s)-T_{\mathrm{pre},i}|\bigr{\}},$
where $T_{\mathrm{pre},i}$ and $T_{\mathrm{obj},i}(s)$ are the efforts recorded immediately before and after lifting the object respectively, $a_{L}$ is a minimum speed factor and $\alpha_{i}\geq 0$ are pre-tuned parameters. State $s$ is the object weight, but is not directly measured.